OpenStudy (anonymous):

Explain how to factor the following trinomials forms: x2 + bx + c and ax2 + bx + c. Is there more than one way to factor this? Show your answer using both words and mathematical notation.

OpenStudy (anonymous):

when we get factoring trinomial we get first x2+bx+c and this are mother flutterers

OpenStudy (ajprincess):

x2 + bx + c let d and e be two numbers such that d*e=c (d and e are factors of c) d+e=b Applying this relation in x2+bx+c u get x2+(d+e)x+d*e =x2+dx+ex+d*e =x(x+d)+e(x+d) =(x+d)(x+e) Upto this part is it clear??